A Bibliography of Publications of Claudio Bartolini

Claudio Bartolini HP Laboratories 1501 Page Mill Rd Palo Alto, CA, 94304 USA

Tel: ?n/a? FAX: ?n/a?

E-mail: claudio.bartolini@hp.com (Internet)

13 October 2017 Version 1.04

Abstract

This bibliography records publications of Claudio Bartolini.

Title word cross-reference

agents [PBB01]. Architecting [BPJ02]. auctions [PBB01]. Automated [BPJ02].

business [TBP02]. business-to-business [TBP02].

commerce [TBP02].

dynamics [PBB01].

e-commerce [TBP02]. Economic [PBB01].

Framework [BPJ02].

lifecycle [TBP02]. multiple [PBB01].

Negotiation [BPJ02].

Reuse [BPJ02].

Semantic [TBP02]. Software [BPJ02].

support [TBP02].

web [TBP02].

References

Bartolini:2002:ARS

[BPJ02] C. Bartolini, C. Preist, and N. R. Jennings. Architecting for reuse: A software framework for automated negotiation. In Third International Workshop on Agent Oriented Software Engineering, 2002. To appear.

Preist:2001:EDA

[PBB01] Chris Preist, Andrew Byde, and Claudio Bartolini. Economic dynamics of agents in multiple auctions. In Proceedings of the fifth international conference on Autonomous agents, pages 545–551. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-326-X.

Trastour:2002:SWS

[TBP02] David Trastour, Claudio Bartolini, and Chris Preist. Semantic web support for the business-to-business ecommerce lifecycle. In Proceedings of the eleventh international conference on World Wide Web, pages 89–98. ACM Press, New York, NY 10036, USA, 2002. ISBN 1-58113-449-5.